how do you solve log base5(x^2+7)=2/3log base5(64)
Can you please show work

Answer by lwsshak3(11628)

how do you solve:
log base5(x^2+7)=2/3log base5(64)
log5(x^2+7)=2/3log5(64)
log5(x^2+7)=log5(64^(2/3))
64^(2/3)=16
log5(x^2+7)=log5(16)
x^2+7=16
x^2=9
x=±3